OpenStudy (anonymous):

5.Find the range of values of p,given that the straight line y=px-1 does not intercept th curve y=x^2+2x+3. @ganeshie8

ganeshie8 (ganeshie8):

yes set both equations equal to each other and get the quadratic in standard form

ganeshie8 (ganeshie8):

\[px - 1 = x^2+2x+3\]

ganeshie8 (ganeshie8):

write it in standard form

ganeshie8 (ganeshie8):

what do you get ?

OpenStudy (anonymous):

\[-x^2-2x+px-3+1=0\]

OpenStudy (anonymous):

\[-x^2-x(2-p)-2=0\]

OpenStudy (anonymous):

\[x^2+x(2-p)+2=0\]

ganeshie8 (ganeshie8):

the constant term doesn't look right, work it again

ganeshie8 (ganeshie8):

\[px - 1 = x^2+2x+3\] subtracting px both sides gives \[- 1 = x^2+x(2-p)+3\] adding 1 both sides gives \[0 = x^2+x(2-p)+4\]

ganeshie8 (ganeshie8):

you want this quadratic equation NOT to have real solutions because you don't want the straight line to touch the curve

ganeshie8 (ganeshie8):

whats the condition for a quadratic NOT to have real solutions ?

OpenStudy (anonymous):

b2-4ac<0

ganeshie8 (ganeshie8):

Yes!

OpenStudy (anonymous):

\[(2-p)^2-4(1)(4)<0\]

ganeshie8 (ganeshie8):

Excellent! see if you can factor the left hand side

OpenStudy (anonymous):

\[4-4p+p^2-16<0\]

OpenStudy (anonymous):

\[p^2-4p-12<0\]

OpenStudy (anonymous):

\[(p-6)(p+2)<0\]

ganeshie8 (ganeshie8):

Looks great! draw a parabola for finding the range

OpenStudy (anonymous):